Nokela Industries purchases a $38.0 million cyclo-converter. The cyclo-converter will be depreciated by $9.5 million per year over four years, starting this year. Suppose Nokelas tax rate is 40%. b. What impact will the cost of the purchase have on the firms cash flow for the next four years? a. What impact will the cost of the purchase have on earnings for each of the next four years? Eamings willbys ilion ach year for four years | by $L1 million each year for four years. (Select from the drop-down menu and round to one decimal p increase decrease
Answer #1

Due to depreciation earnings will be reduced.

It will be subtracted from total profit before calculating tax

Now depreciation is 9.5 million dollar per year and tax Is 40 percent

So tax saved from depretiation = 9.5*0.4 = 3.8 million dollar

So earning will reduce by 9.5-3.8 = 5.7 million dollars

Answer: decrease, $5.7 million per year